OpenStudy (anonymous):

Give the lowest common multiple of 24, 36 & 60

OpenStudy (anonymous):

take the l.c.m

OpenStudy (ash2326):

we have 24, 36, 60 24= 4*3*2 36=4*3*3 60=4*3*5 we have 4 and 3 common in each so that will be the part of LCM 4*3 the uncommon terms are 2 , 3 and 5 so LCM=4*3*2*3*5=360

OpenStudy (anonymous):

24 = 2 x 2 x 2 x 3 36 = 2 x 2 x 3 x 3 60 = 2 x 2 x 3 x 5 lcm = 2 * 2 * 2 * 3 * 3 * 5 = 360

OpenStudy (anonymous):

multiple factors are only included once